### v2

# Rigorous Test for Quantum Integrability and Nonintegrability

Abstract
Whether or not a quantum many-body system is integrable, which is characterized by the presence or absence of local conserved quantities, drastically impacts the dynamics of isolated systems, including thermalization. Nevertheless, a rigorous and comprehensive method for deciding integrability or nonintegrability has remained elusive. In this paper, we address this challenge by introducing rigorously provable tests for integrability and nonintegrability of quantum spin systems with finite-range interactions. Our approach establishes a new paradigm, moving beyond the conventional artisanal methods in the study of nonintegrability. Furthermore, it partially resolves the long-standing conjecture that integrability is governed by the presence or absence of local conserved quantities with small supports. The proposed framework guarantees that the nonintegrability of one-dimensional spin systems with translational symmetry can be confirmed algorithmically, regardless of system size.
I Introduction
A central problem in quantum many-body physics is to understand the equilibrium states and dynamics of strongly interacting quantum many-body systems. Rigorous approaches sometimes provide deep insights into such puzzles. One direction is to establish physically meaningful results for broad classes of systems, such as the Hohenberg–Mermin–Wagner theorem [1, 2] and the Lieb–Schultz–Mattis theorem [3]. Another approach is the use of exactly solvable systems. The Bethe ansatz method, which originates from Bethe’s exact solution of the Heisenberg chain [4], has made it possible to obtain quantities that cannot analytically be calculated for general systems, such as energy eigenvalues, eigenstates, and correlation functions. Furthermore, the algebraic Bethe ansatz method, or the quantum inverse scattering method [5], has revealed that behind such solvability lies integrability, i.e., the existence of an extensive number of local conserved quantities.
However, when considering the dynamics of isolated quantum many-body systems, integrable systems sometimes exhibit anomalous behavior. A typical example appears in thermalization of isolated quantum many-body systems [6]. Recent experiments have shown thermalization, i.e., the relaxation of local observables to thermal equilibrium in well-isolated quantum systems [7]. The eigenstate thermalization hypothesis (ETH) [8, 9, 10], which states that energy eigenstates are locally indistinguishable from thermal equilibrium states, is widely accepted as its mechanism. While the ETH has been numerically demonstrated to hold for various nonintegrable systems [11, 12], it cannot be applied to integrable systems due to the presence of local conserved quantities. Some experiments have also shown that integrable systems exhibit unusual relaxation dynamics [13]. Moreover, the existence of local conserved quantities is known to give rise to nontrivial inequalities in linear response [14, 15]. Thus, significant differences exist between the universal behavior observed in nonintegrable systems and the anomalous behavior in integrable systems. Consequently, in addressing the fundamental question of statistical mechanics—how the empirical laws of macroscopic systems emerge from quantum mechanics—it is crucial to determine whether the system is integrable or nonintegrable.
While integrability has traditionally been a focus of rigorous research, nonintegrability has long remained beyond its scope. However, recent studies have developed methods to prove nonintegrability, or more specifically, the absence of local conserved quantities [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The initial work was limited to a spin-1/2 XYZ-h chain [16], but subsequent studies [17, 21, 22] have classified the nonintegrability of general translationally invariant spin-1/2 chains with nearest-neighbor symmetric interactions. This research has expanded to include spin-1/2 chains with next-nearest-neighbor interactions [18, 20, 27], spin-1 chains [23, 24], and higher-dimensional systems [25, 26]. On the other hand, the existing method is heuristic, relying on a form of “craftsmanship” to individually show the absence of $k$-local conserved quantities for each size $k$ of the operator’s support. Although all the proofs of nonintegrability in previous research have employed similar methods, the underlying unifying structure remains largely unexplored. This lack of a unified framework also poses a significant challenge when determining nonintegrability by using computers because the dimension of the space of $k$-local operators increases exponentially with $k$. Since other numerical tests for nonintegrability, such as the investigation of level spacing statistics [28, 29], are inherently affected by finite-size effects, determining whether the rigorous proof method described above can be formulated as an efficient algorithm is of fundamental importance.
Efficient methods for determining integrability have long been of interest as well in the study of integrable systems [30, 31, 32, 33, 34]. It is conjectured that the existence of $3$-local conserved quantities is a necessary condition for integrability in one-dimensional systems with translational symmetry and nearest-neighbor interactions [34]. Conversely, previous proofs of nonintegrability of such systems [16, 17, 19, 21, 22, 23, 24] have supported the inverse  111 Strictly speaking, it is not literally inverse because there can exist systems that are neither integrable nor nonintegrable, but partially integrable, i.e., with some nontrivial local conserved quantities. Surprisingly, the existence of such partially integrable models is denied in the models examined in previous studies of nonintegrability. conjecture: the absence of $3$-local conserved quantities is a sufficient condition for nonintegrability [21]. However, these tests for integrability and nonintegrability remain empirical, and, to the best of our knowledge, no rigorous proof has yet been established. The main purpose of this paper is to present a rigorously provable test for integrability and nonintegrability based only on $3$-local quantities. This greatly simplifies the existing proof of nonintegrability and gives a partial solution to the above conjectures. Using our rigorous test, we construct an algorithm, which is independent of the system size, to guarantee nonintegrability of translationally invariant systems.
This paper is organized as follows. In Sec. II, we fix the notation and state the main theorem. As an application of the theorem, we prove the absence of local conserved quantities in the XXZ chain with a non-uniform magnetic field. Section III is devoted to the proof of the theorems. Our discussion is based on some graphical language, which is also introduced in this section. Technical details of the proof are provided in Appendices. We also argue that our theorem and its proof can be adopted in the analysis of a related structure: spectrum generating algebras. In Sec. IV, we construct an algorithm of integrability and nonintegrability test for translationally invariant systems. We conclude the paper with a summary and outlook in Sec. V.
II Setup and main result
II.1 Models
We consider a quantum spin chain on $N$ sites subject to the periodic boundary condition. The set of sites is denoted by $\Lambda_{N}=\{1,\dots,N\}$ and we identify $N+i$ with $i$. Each site has a state space $\mathcal{H}\cong\mathbb{C}^{d}$, which is a $d(<\infty)$-dimensional complex Hilbert space. Thus, the whole state space is $\mathcal{H}_{\mathrm{tot}}\coloneqq\mathcal{H}^{\otimes\Lambda_{N}}$. Note that the following argument holds for any finite $d$ and is not limited to spin-1/2 systems.
Next, we introduce $k$-local operators. The set of (bounded) linear operators and that of traceless operators on $\mathcal{H}$ are denoted by $\mathcal{B}$ and $\mathcal{B}_{0}$, respectively. For each operator $\hat{X}$ in $\mathcal{B}_{\mathrm{tot}}\coloneqq\mathcal{B}^{\otimes\Lambda_{N}}$, its support is defined as the smallest set of sites on which $\hat{X}$ acts nontrivially. The set of operators with a support $A(\subset\Lambda_{N})$ is defined as $\mathcal{B}_{0}^{A}\coloneqq\mathcal{B}_{0}^{\otimes A}$. Here we identify $\mathcal{B}^{\otimes A}$ with $\mathcal{B}^{\otimes A}\otimes I^{\otimes(\Lambda_{N}\setminus A)}$ where $I$ is the identity operator. The set of (strictly) $k$-local operators $\mathcal{B}^{(k)}$ is defined as a (direct) sum of $\mathcal{B}_{0}^{A}$ with $\operatorname{\mathrm{size}}(A)=k$, where
$\operatorname{\mathrm{size}}(A)\coloneqq\min\{d\in\mathbb{Z}_{\geq 0}\mid{}^{\exists}l\in\mathbb{Z},\ A+l\subset\{1,\dots,d\}\}$ (1)
is the size of a subset $A$. Note that $\mathcal{B}^{(0)}=\mathcal{B}_{0}^{\emptyset}=\mathbb{C}I^{\otimes\Lambda_{N}}$. We define the set of at most $k$-local quantities ~$\tilde{\mathcal{B}}^{(k)}$ as a (direct) sum of $\mathcal{B}^{(l)}$ for $0\leq l\leq k$. Since $\mathcal{B}=\mathcal{B}_{0}\oplus\mathbb{C}I$ holds, we have ~$\tilde{\mathcal{B}}^{(N)}=\mathcal{B}_{\mathrm{tot}}$. We define a length of an operator $\hat{X}$ as $\operatorname{\mathrm{len}}(\hat{X})\coloneqq\min\{k\in\{0,\dots,N\}\mid\hat{X}\in\tilde{\mathcal{B}}^{(k)}\}$, which represents how local $\hat{X}$ is. When we refer to the operator $\hat{X}$ simply as a $k$-local quantity, we mean $\operatorname{\mathrm{len}}(\hat{X})=k$. Such an operator can be decomposed into $l$-local terms for $0\leq l\leq k$: $\hat{X}=\sum_{l=0}^{k}\hat{X}^{(l)}$ with a nonvanishing $\hat{X}^{(k)}$, where $\hat{X}^{(l)}$ is an element of $\mathcal{B}^{(l)}$. If we emphasize that $\hat{X}^{(l)}$ is an element of $k$-local quantity $\hat{X}$, we write $\hat{X}_{[k]},\hat{X}_{[k]}^{(l)}$ instead of $\hat{X},\hat{X}^{(l)}$. For $l\geq 1$, each $\hat{X}^{(l)}$ is further expanded as $\hat{X}^{(l)}=\sum_{i=1}^{N}\hat{X}^{(l)}_{i}$, where
$\hat{X}^{(l)}_{i}\in\bigoplus_{\begin{subarray}{c}\operatorname{\mathrm{size}}(A)=l\\ i,i+l-1\in A\end{subarray}}\mathcal{B}_{0}^{A}=\mathcal{B}_{0}^{\{i\}}\otimes\mathcal{B}^{\otimes(l-2)}\otimes\mathcal{B}_{0}^{\{i+l-1\}}$ (2)
for $l\geq 2$ and $\hat{X}^{(1)}_{i}\in\mathcal{B}_{0}^{\{i\}}$.
Finally, we introduce a Hamiltonian of this system. We consider nearest-neighbor interactions and on-site potentials. This corresponds to a $2$-local Hamiltonian, which is represented as
$\hat{H}=\sum_{i=1}^{N}\hat{H}^{(2)}_{i}+\sum_{i=1}^{N}\hat{H}^{(1)}_{i},$ (3)
that is, $\hat{H}^{(2)}_{i}\in\mathcal{B}_{0}^{\{i,i+1\}}$ and $\hat{H}^{(1)}_{i}\in\mathcal{B}_{0}^{\{i\}}$. We are mainly interested in translationally invariant models, but we do not assume any relationship between these operators at this point. In addition, $\hat{H}$ can be non-hermitian unless otherwise stated.
II.2 Main Result
We impose the following assumption on $\hat{H}^{(2)}_{i}$:
$\hat{X}\in\mathcal{B}_{0}^{\{i+1\}},\ [\hat{H}^{(2)}_{i},I\otimes\hat{X}]=0$ $\Rightarrow\hat{X}=0$ (4)
$\mbox{and}\ \hat{Y}\in\mathcal{B}_{0}^{\{i\}},\ [\hat{H}^{(2)}_{i},\hat{Y}\otimes I]=0$ $\Rightarrow\hat{Y}=0$
for all $i\in\Lambda_{N}$, which ensures that this Hamiltonian is not “singular”. For example, let us consider an XYZ-type interaction: $\hat{H}^{(2)}_{i}=\sum_{\alpha=x,y,z}J^{\alpha}\hat{S}_{i}^{\alpha}\hat{S}_{i+1}^{\alpha}$ where $\hat{S}_{i}^{\alpha}$ represents the spin operator of spin $\mathcal{S}\in\mathbb{Z}_{>0}/2$. Then, for any $\mathcal{S}$, the assumption (4) is equivalent to the condition that at least two elements in $\{J^{x},J^{y},J^{z}\}$ are nonzero. Otherwise, this interaction becomes an Ising-type one or vanishes.
We are now in a position to state the main result of this paper: a rigorous test for integrability and nonintegrability. Without any further assumptions, we can show the following theorem rigorously.
Theorem 1 (Main Result).
We assume a Hamiltonian $\hat{H}$ with nearest-neighbor interactions and on-site potentials (see Eq. (3)) satisfies the condition (4). If there is no $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$, then the system has no $k$-local conserved quantity for $23\leq k\leq N/$ 2. Equivalently, if there is a $k$-local conserved quantity for $23\leq k\leq N/$ 2, there also exists a $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$.
This theorem provides a sufficient condition for nonintegrability, or more precisely, the absence of $k$-local conserved quantities for $23\leq k\leq N/$ 2. Note that all models satisfying condition (4), among those whose nonintegrability has been proven [16, 19, 21, 22, 23, 24], satisfy this sufficient condition. As an equivalent consequence, Theorem 1 presents the existence of a 3-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$ as a necessary condition for integrability. This yields a partial resolution of the conjecture in Ref. [34], which states that the existence of a 3-local conserved quantity $\hat{Q}$, that is, a 3-local quantity satisfying $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])=0$ is necessary for integrability. In summary, Thm. 1 gives a unified perspective and greatly simplifies existing proofs of nonintegrability, and provides a rigorous foundation for empirical tests of integrability. We note that Thm. 1 can be applied to models with finite-range interactions, spin ladders and models of higher-dimensional systems 222 In this case, the “locality” of an operator is defined in a very weak sense; it is determined by the size of its support for one direction. , because such a model is mathematically equivalent to some spin chain with a larger local dimension.
We can prove a more general theorem, which limits a possible form of conserved quantities. We introduce the following sets of operators:
$\mathcal{B}_{\leq}^{(k)}$$\coloneqq\{\hat{X}\in\mathcal{B}^{(k)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq k\},$ (5)
$\mathcal{B}_{<}^{(k)}$$\coloneqq\{\hat{X}\in\mathcal{B}_{\leq}^{(k)}\oplus\mathcal{B}^{(k-1)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq k-1\}.$ (6)
Note that, since $\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq\operatorname{\mathrm{len}}(\hat{X})+1$ holds, $\hat{X}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$ and $\hat{X}_{[k]}^{(k)}+\hat{X}_{[k]}^{(k-1)}\in\mathcal{B}_{<}^{(k)}$ are the easiest and second easiest conditions that must be satisfied for a $k$-local quantity $\hat{X}_{[k]}=\sum_{l=0}^{k}\hat{X}_{[k]}^{(l)}$ to be a conserved quantity. Under these definitions, the following theorem holds.
Theorem 2.
We have the following isomorphisms between linear spaces:
$\mathcal{B}_{\leq}^{(2)}$ $\cong\mathcal{B}_{\leq}^{(3)}\cong\dots\cong\mathcal{B}_{\leq}^{(N/2)},$ (7)
$\mathcal{B}_{<}^{(3)}$ $\cong\mathcal{B}_{<}^{(4)}\cong\dots\cong\mathcal{B}_{<}^{(N/2)},$ (8)
where the actions of the isomorphisms in Eq. (8) on $\mathcal{B}_{\leq}^{(k)}$ are obtained by the restriction of the domain of the isomorphisms (7).
These isomorphisms yield the stronger result than Thm. 1: if there is no $3$-local quantity $\hat{Q}_{[3]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$, then there is no $k$-local quantity $\hat{Q}_{[k]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[k]},\hat{H}])\leq k-1$ for any $23\leq k\leq N/$ 2. Furthermore, the first isomorphisms (7) provide an insight into how to prove nonintegrability; first, we determine $\mathcal{B}_{\leq}^{(2)}$ and obtain all the elements of $\mathcal{B}_{\leq}^{(3)}$ by Eq. (7). Next, we search for the solution of $\operatorname{\mathrm{len}}([\hat{X}^{(3)}+\hat{X}^{(2)},\hat{H}])\leq 2$ for $\hat{X}^{(3)}\in\mathcal{B}_{\leq}^{(3)}$ and $\hat{X}^{(2)}\in\mathcal{B}^{(2)}\cap(\mathcal{B}^{(2)}_{\leq})^{\perp}$, where we remove the set of trivial solutions $\mathcal{B}^{(2)}_{\leq}\subset\mathcal{B}_{<}^{(3)}$. If there is no such nontrivial solution, we can conclude that the system is nonintegrable. This protocol not only unifies but also greatly simplifies the existing proofs of nonintegrability of systems satisfying (4[16, 19, 21, 22, 23, 24], where the absence of $k$-local conserved quantities is discussed separately for each $k$. A more specific procedure for translationally invariant systems, including how to determine $\mathcal{B}_{\leq}^{(2)}$, will be discussed in Sec. IV.
II.3 Example: nonintegrability of the XXZ chain with a non-uniform magnetic field
As an application of Thm. 1 and Thm. 2, we prove the absence of local conserved quantities in the spin-1/2 XXZ chain with a non-uniform magnetic field. The Hamiltonian of this system is represented as
$\hat{H}=\sum_{i=1}^{N}\sum_{\alpha\in\{x,y,z\}}J_{\alpha}\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{i+1}^{\alpha}+\sum_{i=1}^{N}h_{i}\hat{\sigma}_{i}^{z},$ (9)
where $\hat{\sigma}^{\alpha}$’s represent the Pauli matrices. We assume $J_{x}=J_{y}\neq 0$ and $J_{z}\neq 0$. This is the same form as Eq. (3), where $\hat{H}^{(2)}_{i}$ is the XXZ interaction and $\hat{H}^{(1)}_{i}$ corresponds to a magnetic field. We refer to this magnetic field as uniform when $h_{i}$ takes the same value independently of $i$ ; otherwise, it is called non-uniform. If a magnetic field is uniform, this system is a Bethe solvable system [37], and local conserved quantities are explicitly constructed in Refs. [38, 39].
To consider local conserved quantities in this system, we first determine $\mathcal{B}_{\leq}^{(2)}$. In fact, this step has already been done in Refs. [16, 21, 22], and $\mathcal{B}_{\leq}^{(2)}=\mathbb{C}\hat{H}^{(2)}$ holds. By applying Eq. (7), we can determine $\mathcal{B}_{\leq}^{(k)}$; all of them are one-dimensional. In particular, $\mathcal{B}_{\leq}^{(3)}$ is generated by the following operator:
$\hat{Q}_{[3]}^{(3)}$ $\coloneqq\frac{\mathrm{i}}{2}\sum_{i=1}^{N}[\hat{H}^{(2)}_{i},\hat{H}^{(2)}_{i+1}]$
$=\sum_{i=1}^{N}\sum_{\alpha,\beta,\gamma\in\{x,y,z\}}J_{\alpha}J_{\gamma}\varepsilon_{\alpha\beta\gamma}\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{i+1}^{\beta}\hat{\sigma}_{i+2}^{\gamma}.$ (10)
Here, $\varepsilon_{\alpha\beta\gamma}$ denotes the Levi-Civita symbol.
Next, we determine $\mathcal{B}_{<}^{(3)}$. In the following, we show that this set is trivial, that is, $\mathcal{B}_{<}^{(3)}\subset\mathcal{B}^{(2)}$ if a magnetic field is non-uniform. To prove this, we derive a contradiction by assuming that
$\operatorname{\mathrm{len}}([\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)},\hat{H}])\leq 2$ (11)
holds for some $\hat{Q}_{[3]}^{(2)}\in\mathcal{B}^{(2)}$. This operator can be expanded in terms of the Pauli operators as
$\hat{Q}_{[3]}^{(2)}=\sum_{i=1}^{N}\sum_{\alpha,\beta\in\{x,y,z\}}q_{\alpha\beta,i}\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{i+1}^{\beta}.$ (12)
Now we expand $[\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)},\hat{H}]$ as
$[\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)},\hat{H}]$ $=2\mathrm{i}\sum_{i=1}^{N}\sum_{\alpha,\beta,\gamma\in\{x,y,z\}}c_{\alpha\beta\gamma,i}\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{i+1}^{\beta}\hat{\sigma}_{i+2}^{\gamma}$
[…]
where $c_{\alpha\beta\gamma,i}$ is a function of $J_{\alpha^{\prime}}$ and $q_{\alpha^{\prime}\beta^{\prime},i}$. The assumption (11) is equivalent to $c_{\alpha\beta\gamma,i}=0$ for all $\alpha,\beta,\gamma$ and $i\in\Lambda_{N}$. First, we calculate $c_{xxz,i}$. This coefficient comes from the following three commutators:
$[J_{z}J_{x}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{y}\hat{\sigma}_{i+2}^{z},h_{i+1}\hat{\sigma}_{i+1}^{z}],$ (14)
$[-J_{z}J_{x}\hat{\sigma}_{i}^{y}\hat{\sigma}_{i+1}^{x}\hat{\sigma}_{i+2}^{z},h_{i}\hat{\sigma}_{i}^{z}],\ \mbox{and}$ (15)
$[q_{xy,i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{y},J_{z}\hat{\sigma}_{i+1}^{z}\hat{\sigma}_{i+2}^{z}],$ (16)
which determine the form of $c_{xxz,i}$:
$c_{xxz,i}=J_{z}[J_{x}(h_{i+1}-h_{i})+q_{xy,i}].$ (17)
Next, we calculate $c_{zyy,i-1}$. This coefficient comes from the following three commutators:
$[J_{z}J_{x}\hat{\sigma}_{i-1}^{z}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{y},h_{i}\hat{\sigma}_{i}^{z}],$ (18)
$[-J_{z}J_{x}\hat{\sigma}_{i-1}^{z}\hat{\sigma}_{i}^{y}\hat{\sigma}_{i+1}^{x},h_{i+1}\hat{\sigma}_{i+1}^{z}],\ \mbox{and}$ (19)
$[q_{xy,i}\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+1}^{y},J_{z}\hat{\sigma}_{i-1}^{z}\hat{\sigma}_{i}^{z}],$ (20)
which determine the form of $c_{zyy,i-1}$:
$c_{zyy,i-1}=J_{z}[J_{x}(h_{i+1}-h_{i})-q_{xy,i}].$ (21)
Therefore, $c_{xxz,i}=c_{zyy,i-1}=0$ implies $h_{i+1}=h_{i}$, where we use $J_{x},J_{z}\neq 0$. This contradicts with non-uniformity of a magnetic field, and we have $\mathcal{B}_{<}^{(3)}\subset\mathcal{B}^{(2)}$. By applying Thm. 2, it turns out that $\mathcal{B}_{<}^{(k)}$ is also trivial for every $23\leq k\leq N/$ 2.
In summary, we have proven the following result.
Proposition 3.
Any model in Eq. (9) with a non-uniform magnetic field has no $k$-local conserved quantity for $23\leq k\leq N/$ 2.
Although nonintegrability of systems without spatial uniformity is also discussed in Ref. [18], the above result provides the first example in which it is rigorously shown that the system is nonintegrable only if it is not uniform.
We comment on the connection between the above result and thermalization of this system. Many-body localization (MBL) is a phenomenon where strong disorder prevents a system from thermalizing, leading to a breakdown of the eigenstate thermalization hypothesis (ETH)  [40, 41, 42, 43]. Whether the XXZ model with a random magnetic field exhibits MBL in the thermodynamic limit is a subject of debate [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. The MBL is attributed to the existence of an extensive number of quasi-local conserved quantities [44, 54], which are operators consisting of $k$-local operators ($0\leq k\leq N$) whose coefficients decay exponentially with $k$. Such a mechanism is sometimes referred to as the emergence of local integrability [42, 43], but note that this is different from the integrability we discuss, in the sense that there is an extensive number of (strictly) local conserved quantities. In fact, Prop. 3 tells us that the XXZ model with a random magnetic field is nonintegrable in our sense.
III Outline of Proof
III.1 String Diagram
Before proving Thm. 2, we introduce a “basis-independent” approach for a proof of nonintegrability. We adopt a string diagram as a useful graphical language, which is widely used in the various fields of physics and mathematics [55, 56, 57]. The basic idea is to represent a linear map by a box and wires. For example, $\hat{X}\in\mathcal{B}\otimes\mathcal{B}_{0}$, $f:\mathcal{B}_{0}^{\otimes 3}\to\mathcal{B}_{0}^{\otimes 2}$, and $(\mathrm{id}_{\mathcal{B}}\otimes f)(\hat{X}\otimes\bullet):\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}\otimes\mathcal{B}_{0}^{\otimes 2}$ are denoted as
[…]
respectively, where $\mathrm{id}_{\mathcal{B}}$ is the identity map on $\mathcal{B}$. Here, each box represents a linear map and the top line corresponds to an input and the bottom line corresponds to an output. In other words, time flows from top to bottom. Note that an element of a linear space is identified with a linear map from $\mathbb{C}$, which is denoted as the “empty” line. The horizontal direction corresponds to the spatial direction, and placing the boxes horizontally corresponds to their tensor product. In the following, solid and dashed lines represent $\mathcal{B}_{0}$ and $\mathcal{B}$, respectively. Using this expression, the injectivity of $[\hat{H}^{(2)}_{i},\bullet\otimes I]$ and $[\hat{H}^{(2)}_{i},I\otimes\bullet]$ (see Eq. (4)) is equivalent to the existence of linear maps $\mu_{i},\lambda_{i}:\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}_{0}$ satisfying
[…]
for all $i\in\Lambda_{N}$. Here, $[\ ,\ ]:\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}_{0}$ denotes the commutator on $\mathcal{B}_{0}$.
III.2 Isomorphisms between $\mathcal{B}^{(k)}_{\leq}$’s
Now we move on to the proof of Eq. (7). It is sufficient to show the following proposition.
Proposition 4.
We fix an integer $k$ such that $23\leq k\leq N/$ 2. For each $\hat{Q}_{[k]}^{(k)}=\sum_{i=1}^{N}\hat{Q}^{(k)}_{[k],i}\in\mathcal{B}_{\leq}^{(k)}$, there (uniquely) exists $\hat{Q}_{[k-1]}^{(k-1)}=\sum_{i=1}^{N}\hat{Q}^{(k-1)}_{[k-1],i}\in\mathcal{B}_{\leq}^{(k-1)}$ satisfying
[…]
Proof.
The assumption $\operatorname{\mathrm{len}}([\hat{Q}_{[k]}^{(k)},\hat{H}])\leq k$ is equivalent to the following equations:
[Uncaptioned image] (25)
for all $i\in\Lambda_{N}$. This equation shows that the second wire from the left in $\hat{Q}^{(k)}_{[k],i}$ is, in fact, a solid line. Furthermore, by using the injectivity (23), we find that Eq. (24) is satisfied by taking $\hat{Q}^{(k-1)}_{[k-1],i}$ as
[…]
By substituting Eq. (24) into Eq. (25) and using the injectivity (23), we obtain
[Uncaptioned image] (27)
for all $i\in\Lambda_{N}$, which is equivalent to the condition $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$. The uniqueness of $\hat{Q}_{[k-1]}^{(k-1)}$ is also shown by the injectivity (23). ∎
If we define $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}^{(k)}$ as in Eq. (24) for some $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$, then it is easy to verify $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$. Therefore, we have shown $\mathcal{B}_{\leq}^{(k)}\cong\mathcal{B}_{\leq}^{(k-1)}$ and Eq. (7). By applying Eq. (24) iteratively, $\hat{Q}^{(k)}_{[k],i}$ is represented as
[…]
where $\hat{Q}_{[2]}^{(2)}=\sum_{i=1}^{N}\hat{H}^{(2)}_{[2],i}$ is an element of $\mathcal{B}^{(2)}_{\leq}$. We note that, if $\hat{Q}_{[2]}^{(2)}=\hat{H}^{(2)}$, this operator has the same form as the $3$-local conserved quantities of Bethe solvable fundamental systems [34] (see Eq. (10)).
III.3 Isomorphisms between <$\mathcal{B}^{(k)}_{<}$’s
We conclude this section by proving Eq. (8). Here, we shall prove only the case of $\mathcal{B}_{<}^{(3)}\cong\mathcal{B}_{<}^{(4)}$ to explain the bare essentials of the proof. We give a mathematically rigorous and complete proof of Eq. (8) in Appx. B for the interested readers.
We introduce a new symbol to represent the commutator between $2$-local operators. We denote a linear map $[\bullet,\hat{X}]:\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}^{\otimes 2}$ as
[…]
If one line of the output is represented by a solid line instead of a dashed one, it indicates that a projection:
$\mathcal{B}\ni\hat{A}\mapsto\hat{A}-\frac{\operatorname{\mathrm{Tr}}\hat{A}}{d}I\in\mathcal{B}_{0}$ (30)
from $\mathcal{B}$ to $\mathcal{B}_{0}$ is performed at the corresponding site.
For $\alpha$ defined above, the, the following lemma holds, which is essential for the proof of Eq. (8).
Lemma 5.
For $\hat{C}\in\mathcal{B}_{0}^{\otimes 2}$, the following identity holds:
[…]
We provide a proof of this lemma in Appx. A. By using this lemma, we can show the following proposition.
Proposition 6.
There is an isomorphism
$\mathcal{B}_{<}^{(4)}\ni\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\mapsto\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<}^{(3)},$ (32)
where $\hat{Q}_{[4]}^{(4)}$ and $\hat{Q}_{[3]}^{(3)}$ satisfy Eq. (24) for $k=4$.
Proof.
First we construct a map from $\mathcal{B}_{<}^{(4)}$ to $\mathcal{B}_{<}^{(3)}$. For $\hat{Q}_{[4]}^{(4)}\in\mathcal{B}_{\leq}^{(4)}$ and $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<}^{(4)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$, where $\hat{Q}_{[2]}^{(2)}$ is defined by Eq. (28). Here, $[\ ,\hat{X}]:\mathcal{B}_{0}\ni\hat{Y}\mapsto[\hat{Y},\hat{X}]\in\mathcal{B}_{0}$ is the commutator with $\hat{X}\in\mathcal{B}_{0}$. Let us focus on the terms in the first bracket on the right-hand side (RHS) of Eq. (33). By applying Lem. 5 to the middle term, we can decompose these terms into the following form:
[…]
where $\hat{A}^{(3)}_{L}$ and $\hat{A}^{(3)}_{R}$ satisfy
[…]
Similarly, by using the Jacobi identity, the terms in the second bracket on the RHS of Eq. (33) is decomposed into a similar form as Eq. (34333 See Appx. B for details. , by two operators $\hat{B}^{(3)}_{L}$ and $\hat{B}^{(3)}_{R}$ satisfying
$\hat{B}^{(3)}_{L,i}+\hat{B}^{(3)}_{R,i}$
[…]
The above relations allow us to simplify Eq. (33) as
[…]
where two operators $\hat{Q^{\prime}}^{(3)}_{L,i}$ and $\hat{Q^{\prime}}^{(3)}_{R,i}$ are defined as
$\hat{Q^{\prime}}^{(3)}_{L,i}$ $\coloneqq\hat{Q}^{(3)}_{[4],i}+\hat{A}^{(3)}_{L,i}+\hat{B}^{(3)}_{L,i},$ (38)
$\hat{Q^{\prime}}^{(3)}_{R,i}$ $\coloneqq\hat{Q}^{(3)}_{[4],i}-\hat{A}^{(3)}_{R,i}-\hat{B}^{(3)}_{R,i}.$ (39)
Similarly to Prop. 4, there is an operator $\hat{Q^{\prime}}^{(2)}_{i+1}\in\mathcal{B}_{0}^{\otimes 2}$ satisfying
[…]
Equations (38) and (39) yield the following equality:
$\hat{Q^{\prime}}^{(3)}_{L,i}-\hat{Q^{\prime}}^{(3)}_{R,i}=\hat{A}^{(3)}_{L,i}+\hat{B}^{(3)}_{L,i}+\hat{A}^{(3)}_{R,i}+\hat{B}^{(3)}_{R,i}.$ (42)
By substituting Eqs. (35), (36), (40) and (41) into Eq. (42), we obtain
[Uncaptioned image] (43)
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<}^{(3)}$, where we define $\hat{Q}_{[3]}^{(2)}$ as $\frac{2}{3}\sum_{i=1}^{N}\hat{Q^{\prime}}^{(2)}_{i}$. Therefore, we have constructed a linear transformation $i_{34}:\mathcal{B}_{<}^{(4)}\to\mathcal{B}_{<}^{(3)}$. The injectivity of this map is confirmed from the injectivity (23). Conversely, by using the above equalities (e.g., Eq. (38)), we can construct a mapping $i_{43}:\mathcal{B}_{<}^{(3)}\to\mathcal{B}_{<}^{(4)}$ satisfying $i_{34}\circ i_{43}=\mathrm{id}_{\mathcal{B}_{<}^{(3)}}$. This implies the surjectivity of $i_{34}$, which completes the proof. ∎
III.4 Absence of spectrum generating algebras
As discussed in Ref. [25], the proof of the absence of local conserved quantities is applicable with a slight modification to the analysis of a related algebraic structure: spectrum generating algebras. A system with a Hamiltonian $\hat{H}$ is said to admit a spectrum generating algebra (SGA) for an operator $\hat{Q}$ if
$[\hat{Q},\hat{H}]=\mathcal{E}\hat{Q}$ (44)
holds for some $\mathcal{E}\in\mathbb{R}\setminus\{0\}$. This algebraic structure is a variant of the condition $[\hat{Q},\hat{H}]=0$ for the existense of a conserved quantity, and $\eta$-pairing [59, 60] in the Hubbard model is a well-known example.
In parallel with Thm. 1, the following theorem holds.
Theorem 7.
Let us consider a Hamiltonian $\hat{H}$ of the same form as in Thm. 1. If $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}]-\mathcal{E}\hat{Q})\leq 2$ does not hold for any $\mathcal{E}\in\mathbb{R}\setminus\{0\}$ and any $3$-local quantity $\hat{Q}$, then the system has no $k$-local quantity satisfying Eq. (44) for some $\mathcal{E}\in\mathbb{R}\setminus\{0\}$ and $23\leq k\leq N/$ 2. In other words, the system does not admit an SGA for a $k$-local quantity.
Proof.
As a generalization of $\mathcal{B}_{<}^{(k)}$, we introduce the following set of operators:
$\mathcal{B}_{<,\mathcal{E}}^{(k)}\coloneqq\{\hat{X}\in\mathcal{B}^{(k)}\oplus\mathcal{B}^{(k-1)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}]-\mathcal{E}\hat{X})\leq k-1\}.$ (45)
Note that $\mathcal{B}_{<,\mathcal{E}}^{(k)}\subset\mathcal{B}_{\leq}^{(k)}\oplus\mathcal{B}^{(k-1)}$ holds. To prove Thm. 7 for the case of $k=4$, we first construct a linear transformation ${i^{\prime}}_{34}:\mathcal{B}_{<,\mathcal{E}}^{(4)}\to\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
In fact, this map can be defined in a manner similar to ${i}_{34}:\mathcal{B}_{<}^{(4)}\to\mathcal{B}_{<}^{(3)}$. For $\hat{Q}_{[4]}^{(4)}\in\mathcal{B}_{\leq}^{(4)}$ and $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<,\mathcal{E}}^{(4)}$ is equivalent to the following equalities:
[…]
for all $i\in\Lambda_{N}$. As in the proof of the previous subsection, Eq. (46) is equivalent to Eq. (37), where $\hat{Q^{\prime}}^{(3)}_{L,i}$ is redefined as
$\hat{Q^{\prime}}^{(3)}_{L,i}\coloneqq\hat{Q}^{(3)}_{[4],i}+\hat{A}^{(3)}_{L,i}+\hat{B}^{(3)}_{L,i}-\mathcal{E}\hat{Q}^{(3)}_{[3],i}.$ (47)
Under this definition, we can construct $\hat{Q}_{[3]}^{(2)}\in\mathcal{B}^{(2)}$ satisfying
[…]
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
On the other hand, for any $\mathcal{E}\neq 0$, $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$ implies $\hat{Q}_{[3]}^{(3)}=0$ by the assumption. Therefore, we have $\hat{Q}_{[4]}^{(4)}=0$, or equivalently, $\mathcal{B}_{<,\mathcal{E}}^{(4)}\subset\mathcal{B}^{(3)}$ for every $\mathcal{E}\in\mathbb{R}\setminus\{0\}$, which indicates that there is no $4$-local quantity satisfying Eq. (44). With a slight modification of the proof of Thm. 2 for the general $k$ case (Appx. B), we can show the general $k$ case as well. ∎
We note that SGAs are of interest in the study of thermalization. If an operator $\hat{Q}$ satisfies Eq. (44) on some subspaces, the system is said to admit a restricted spectrum generating algebra [61, 62]. This structure provides a unified origin of the tower of quantum many-body scar states, which prevents thermalization.
IV Algorithm of integrability test
By using Thm. 1 and Thm. 2, we can construct an algorithm of test for integrability and nonintegrability. In this section, we only consider translationally invariant Hamiltonians, for simplicity. In this case, we can decompose the set of $k$-local operators into momentum sectors by the Fourier transformation as
$\mathcal{B}^{(k)}=\bigoplus_{p\in\bigl{\{}e^{\frac{2\pi\mathrm{i}}{N}m}\bigm{|}m=0,\dots,N-1\bigr{\}}}\mathcal{B}^{(k,p)},$ (49)
where $p$ represents the momentum, which is the eigenvalue of the translation operation $\tau:\mathcal{B}^{\otimes A}\to\mathcal{B}^{\otimes(A-1)}$. Since this decomposition is consistent with the subspaces $\mathcal{B}^{(k)}_{\leq}$ and <$\mathcal{B}^{(k)}_{<}$, the integrability test introduced in Sec. II.2 is reduced to the following algorithm. First, we determine the space $\mathcal{B}^{(2,p)}_{\leq}\coloneqq\mathcal{B}^{(2)}_{\leq}\cap\mathcal{B}^{(2,p)}$. This step is equivalent to the following generalized eigenvalue problem:
[Uncaptioned image] (50)
for $\hat{Q}_{[2],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$, where we use the translational invariance $\tau(H^{(2)}_{i+1})=H^{(2)}_{i}$ and the fact that $\hat{Q}_{[2]}^{(2)}\in\mathcal{B}^{(2,p)}$ satisfies $\tau(\hat{Q}^{(2)}_{[2],i+1})=p\hat{Q}^{(2)}_{[2],i}$. We note that $\mathcal{B}^{(2,p)}_{\leq}=\{0\}$ except for at most $(d^{2}-1)^{2}$ numbers of ($N$-independent) momentum $p$ by the injectivity (23).
Next, we check whether <$\mathcal{B}^{(3)}_{<}$ is trivial (i.e., equal to $\mathcal{B}^{(2)}_{\leq}$) or not. This step is equivalent to finding solutions of the following equality:
[…]
for $\hat{Q}_{[2],1}^{(2)}$ satisfying Eq. (50) and for $\hat{Q}_{[3],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$. If the above equality implies $\hat{Q}_{[2],1}^{(2)}=0$, we can conclude that this system is nonintegrable by Thm. 1. The above procedure consists of eigenvalue problems for a finite number of linear maps, which are independent of the system size.
As pointed out below Thm. 1, this algorithm can also be applied to a system of spin ladders and models with finite-range interactions. As discussed in Refs. [25, 26], the proof of nonintegrability of systems on hypercubic lattices with nearest-neighbor interactions can be reduced to that of ladder systems. Hence, we can indirectly deal with higher-dimensional systems as well.
Note that given additional pieces of information about the Hamiltonian, it may be possible to construct a less costly algorithm. For example, suppose that there exists a subspace $\mathcal{A}$ of $\mathcal{B}_{0}$ satisfying $\hat{H}_{1}^{(1)}\in\mathcal{A}$ and $\hat{H}_{1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Then, by Eq. (50), we have $\hat{Q}_{[2],1}^{(2)}\in\mathcal{A}^{\otimes 2}$. Furthermore, a solution of Eq. (51) satisfies $\hat{Q}_{[3],1}^{(2)}\in\mathcal{A}_{2}^{\otimes 2}$, where $\mathcal{A}_{2}$ is defined as
$\mathcal{A}_{2}\coloneqq\operatorname{\mathrm{Span}}\{\hat{A}\hat{B}\mid\hat{A},\hat{B}\in\mathcal{A}\}\cap\mathcal{B}_{0}.$ (52)
This observation is useful to show the nonintegrability of general spin-$S$ systems. As a simple example, let us consider the spin-$S$ Heisenberg chain. In this case, $\mathcal{A}$ consists of a linear combination of spin operators $\hat{S}^{\alpha}$, and $\mathcal{A}_{2}$ is $8$-dimensional regardless of $S$ if $S\geq 1$. Therefore, we can show the nonintegrability of the Heisenberg chain for $S\geq 3/2$ as well as for $S=1$ [23, 24].
V Conclusion and outlook
In this work, we have proven a sufficient condition using only 3-local quantities that guarantees nonintegrability, i.e., the absence of local conserved quantities. This result reveals a unified structure behind the existing proofs of nonintegrability, independent of the specific system, and provides a partial solution to the long-standing conjecture that the existence of 3-local conserved quantities is a necessary condition for integrability. Using the obtained theorem, we have also constructed a procedure for determining whether a given translationally invariant system is integrable or nonintegrable, independent of system size. This opens the way for numerical determination of nonintegrability with a rigorous basis.
One straightforward extension of this work is a unified understanding of the proof of nonintegrability in systems that do not satisfy the assumption (4). Typical examples of such systems are those with Ising-type interactions, which are investigated in Refs. [17, 26]. In these proofs of nonintegrability, it may be necessary to consider up to $(k-1)$-local outputs to show the absence of $k$-local conserved quantities [17]. Thus, the structure of nonintegrability may be qualitatively different from the nonintegrability of the system satisfying the assumption (4).
Another challenging direction is to discuss the existence or absence of structures other than local conserved quantities. In Sec. II.3, we have demonstrated that even in the MBL phase, there is no (strictly) local conserved quantity. In the context of distinguing the MBL phase and the ergodic phase where the ETH holds, a crucial question is whether the method of proofs of nonintegrability can be applied to quasi-local conserved quantities. As a related subject, we have discussed the absence of SGAs in Sec. III.4. If RSGAs could be analyzed, it would be possible to determine the presence or absence of quantum many-body scarring. These generalizations can lead to further classification of many-body systems beyond the presence or absence of local conserved quantities.
Acknowledgements.
The author would like to thank Hosho Katsura, Mizuki Yamaguchi, Yuya Chiba, and Naoto Shiraishi for fruitful discuccions, and Masahito Ueda, Masaya Nakagawa and Hal Tasaki for their valuable comments on the manuscript. The author is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), World-Leading INnovative Graduate Study (WINGS) Program, The University of Tokyo.
Data availability
No datasets were generated or analyzed during the current study.
Appendix A Proof of Lemma 5
We expand an operator $\hat{C}\in\mathcal{B}_{0}^{\otimes 2}$ as $\hat{C}=\sum_{a}\hat{X}_{a}\otimes\hat{Y}_{a}$. Then, the RHS of Eq. (31) is expanded as
$\hat{D}\otimes\hat{E}$
$\mapsto$ $\sum_{a,b}[[[\hat{D},\hat{X}_{a}]\otimes\hat{Y}_{a},\hat{X}_{b}\otimes\hat{Y}_{b}],I\otimes\hat{E}]$
$\ +[\hat{D}\otimes I,[\hat{X}_{b}\otimes[\hat{Y}_{b},\hat{E}],\hat{X}_{a}\otimes\hat{Y}_{a}]]$
$=$ $\sum_{a,b}[\hat{D},\hat{X}_{a}]\hat{X}_{b}\otimes[\hat{Y}_{a}\hat{Y}_{b},\hat{E}]-\hat{X}_{b}[\hat{D},\hat{X}_{a}]\otimes[\hat{Y}_{b}\hat{Y}_{a},\hat{E}]$
$\ +[\hat{D},\hat{X}_{b}\hat{X}_{a}]\otimes[\hat{Y}_{b},\hat{E}]\hat{Y}_{a}-[\hat{D},\hat{X}_{a}\hat{X}_{b}]\otimes\hat{Y}_{a}[\hat{Y}_{b},\hat{E}]$
$=$ $\sum_{a,b}[\hat{D},\hat{X}_{a}]\hat{X}_{b}\otimes[\hat{Y}_{a},\hat{E}]\hat{Y}_{b}-\hat{X}_{b}[\hat{D},\hat{X}_{a}]\otimes\hat{Y}_{b}[\hat{Y}_{a},\hat{E}]$
$\ +[\hat{D},\hat{X}_{b}]\hat{X}_{a}\otimes[\hat{Y}_{b},\hat{E}]\hat{Y}_{a}-\hat{X}_{a}[\hat{D},\hat{X}_{b}]\otimes\hat{Y}_{a}[\hat{Y}_{b},\hat{E}]$
$=$ $2\sum_{a,b}[[\hat{D},\hat{X}_{a}]\otimes[\hat{Y}_{a},\hat{E}],\hat{X}_{b}\otimes\hat{Y}_{b}],$ (53)
where we use the Leibniz rule of commutator: $[AB,C]=[A,C]B+A[B,C]$ in the second equality. The last line in Eq. (53) corresponds to the left-hand side of Eq. (31). $\square$
Appendix B Complete Proof of Theorem 2
We provide complete the proof of Eq. (8). It is sufficient to show the following proposition.
Proposition 8.
Let $k\geq 4$. There is an isomorphism
$\mathcal{B}_{<}^{(k)}\ni\hat{Q}_{[k]}^{(k)}+\hat{Q}_{[k]}^{(k-1)}\mapsto\hat{Q}_{[k-1]}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<}^{(k-1)},$ (54)
where $\hat{Q}_{[k]}^{(k)}$ and $\hat{Q}_{[k-1]}^{(k-1)}$ satisfy Eq. (24).
The proof of Prop. 8 follows the same procedure as for Prop. 6. First we construct a map from $\mathcal{B}_{<}^{(k)}$ to $\mathcal{B}_{<}^{(k-1)}$. For $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$ and $\hat{Q}_{[k]}^{(k-1)}\in\mathcal{B}^{(k-1)}$, the condition $\hat{Q}^{(k)}+\hat{Q^{\prime}}^{(k-1)}\in\mathcal{B}_{<}^{(k)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$, where $\hat{Q}_{[2]}^{(2)}$ is defined by Eq. (28). We denote the terms in the first and second brackets on the RHS of Eq. (55) as $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$, respectively. Similarly to the case of $k=4$, we can decompose $\hat{F}^{(k)}_{i}$ and $\hat{G}^{(k)}_{i}$ in terms of operators $\hat{A}^{(k-1)}_{L},\hat{A}^{(k-1)}_{R},\hat{B}^{(k-1)}_{L}$, and $\hat{B}^{(k-1)}_{R}$. Specifically, we have the following lemma for $\hat{F}^{(k)}_{i}$.
Lemma 9.
There exist operators $\hat{A}^{(k-1)}_{L},\hat{A}^{(k-1)}_{R}\in\mathcal{B}^{(k-1)}$ satisfying
[…]
$\hat{F}^{(k-1)}_{i}$ $=\frac{k-2}{k-1}\left(\hat{A}^{(k-1)}_{L,i}+\hat{A}^{(k-1)}_{R,i}\right).$ (57)
Similarly, the following lemma holds for $\hat{G}^{(k)}_{i}$.
Lemma 10.
There exist operators $\hat{B}^{(k-1)}_{L},\hat{B}^{(k-1)}_{R}\in\mathcal{B}^{(k-1)}$ satisfying
[…]
$\hat{G}^{(k-1)}_{i}$ $=\frac{k-2}{k-1}\left(\hat{B}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{R,i}\right).$ (59)
By combining the above two lemmas, we can show Prop. 8.
Proof of Prop. 8.
We define the following two operators:
$\hat{Q^{\prime}}^{(k-1)}_{L,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}+\hat{A}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{L,i},$ (60)
$\hat{Q^{\prime}}^{(k-1)}_{R,i}$ $\coloneqq\hat{Q}^{(k-1)}_{[k],i}-\hat{A}^{(k-1)}_{R,i}-\hat{B}^{(k-1)}_{R,i}.$ (61)
By definition, we have
$\hat{Q^{\prime}}^{(k-1)}_{L,i}-\hat{Q^{\prime}}^{(k-1)}_{R,i}=\hat{A}^{(k-1)}_{L,i}+\hat{A}^{(k-1)}_{R,i}+\hat{B}^{(k-1)}_{L,i}+\hat{B}^{(k-1)}_{R,i}.$ (62)
Furthermore, by using Eq. (55), we can find an operator $\hat{Q^{\prime}}^{(k-2)}_{i+1}\in\mathcal{B}_{0}^{\otimes(k-2)}$ satisfying
[…]
Therefore, for $\hat{Q}_{[k-1]}^{(k-2)}\coloneqq\frac{k-2}{k-1}\sum_{i}\hat{Q^{\prime}}^{(k-2)}_{i}$ we have
[…]
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[k-1]}^{(k-1)}+\hat{Q}_{[k-1]}^{(k-2)}\in\mathcal{B}_{<}^{(k-1)}$. A linear map from $\mathcal{B}_{<}^{(k)}$ to $\mathcal{B}_{<}^{(k-1)}$ is now constructed. By the same discussion as in the main text, this is in fact an isomorphism. ∎
The remaining task is to prove Lem. 9 and Lem. 10.
Proof of Lem. 9.
The operator $\hat{F}^{(k)}_{i}$ consists of $k-1$ operators $\hat{F}^{(k)}_{i,0},\dots,\hat{F}^{(k)}_{i,k-2}$, where each $\hat{F}^{(k)}_{i,l}$ is defined as
$\hat{F}^{(k)}_{i,l}\coloneqq[\hat{Q}^{(k)}_{[k],i},\hat{H}^{(2)}_{i+l}].$ (66)
By using Lem. 5, we can decompose $\hat{F}^{(k)}_{i,l}$ for $31\leq l\leq k-$ 3 as
[…]
$\eqqcolon\frac{1}{2}(\hat{F}^{(k)}_{L,i,l}+\hat{F}^{(k)}_{R,i,l}).$ (67)
For $l=0$ and $l=k-2$, we define $\hat{F}^{(k)}_{L/R,i,l}$ as
$\hat{F}^{(k)}_{L,i,l}=\hat{F}^{(k)}_{R,i,l}=\hat{F}^{(k)}_{i,l}.$ (68)
Note that $\hat{F}^{(k)}_{L,i,l}$ and $\hat{F}^{(k)}_{R,i,l}$ satisfy the following equalities:
[…]
In addition, for $l\geq 2$ we have
[…]
Similarly, for $l\leq k-4$ we have
[…]
We now construct operators $\hat{A}^{(k-1)}_{L},\hat{A}^{(k-1)}_{R}\in\mathcal{B}^{(k-1)}$ as
$\hat{A}^{(k-1)}_{L,i}$ $\coloneqq\frac{1}{k-2}\sum_{l=0}^{k-3}\left(\frac{l+2}{2}\hat{F}^{(k-1)}_{L,i,k-3-l}+\frac{l}{2}\hat{F}^{(k-1)}_{R,i,k-3-l}\right),$ (73)
$\hat{A}^{(k-1)}_{R,i}$ $\coloneqq\frac{1}{k-2}\sum_{l=0}^{k-3}\left(\frac{l}{2}\hat{F}^{(k-1)}_{L,i,l}+\frac{l+2}{2}\hat{F}^{(k-1)}_{R,i,l}\right).$ (74)
By using Eqs. (69), (70), (71), and (72), we can show that $\hat{A}^{(k-1)}_{L}$ and $\hat{A}^{(k-1)}_{R}$ satisfy Eq. (56). Furthermore, Eq. (57) follows from Eq. (67), which completes this proof. ∎
Proof of Lem. 10.
The operator $\hat{G}^{(k)}_{i}$ consists of $k$ operators $\hat{G}^{(k)}_{i,0},\dots,\hat{G}^{(k)}_{i,k-1}$, where each $\hat{G}^{(k)}_{i,l}$ is defined as
$\hat{G}^{(k)}_{i,l}\coloneqq[\hat{Q}^{(k)}_{[k],i},\hat{H}^{(1)}_{i+l}].$ (75)
We can decompose $\hat{G}^{(k)}_{i,l}$ for $21\leq l\leq k-$ 2 as
[…]
$\eqqcolon\hat{G}^{(k)}_{L,i,l}+\hat{G}^{(k)}_{R,i,l},$ (76)
where we use the Jacobi identity: $[[\hat{A},\hat{B}],\hat{C}]=[[\hat{A},\hat{C}],\hat{B}]+[\hat{A},[\hat{B},\hat{C}]]$ in the second equality. For $l=0$ and $l=k-2$, we define $\hat{G}^{(k)}_{L/R,i,l}$ as
$\hat{G}^{(k)}_{R,i,0}$ $=\hat{G}^{(k)}_{i,0},$ (77)
$\hat{G}^{(k)}_{L,i,k-1}$ $=\hat{G}^{(k)}_{i,k-1},$ (78)
$\hat{G}^{(k)}_{L,i,0}$ $=\hat{G}^{(k)}_{R,i,k-1}=0.$ (79)
Note that $\hat{G}^{(k)}_{L,i,l}$ and $\hat{G}^{(k)}_{R,i,l}$ satisfy the following equalities:
[…]
for $21\leq l\leq k-$ 2. In addition, for $l\geq 2$ we have
[…]
Similarly, for $l\leq k-3$ we have
[…]
We now construct operators $\hat{B}^{(k-1)}_{L},\hat{B}^{(k-1)}_{R}\in\mathcal{B}^{(k-1)}$ as
$\hat{B}^{(k-1)}_{L,i}$ $\coloneqq\frac{1}{k-2}\sum_{l=0}^{k-2}\left((l+1)\hat{G}^{(k-1)}_{L,i,k-2-l}+l\hat{G}^{(k-1)}_{R,i,k-2-l}\right),$ (84)
$\hat{B}^{(k-1)}_{R,i}$ $\coloneqq\frac{1}{k-2}\sum_{l=0}^{k-2}\left(l\hat{G}^{(k-1)}_{L,i,l}+(l+1)\hat{G}^{(k-1)}_{R,i,l}\right).$ (85)
By using Eqs. (80), (81), (82), and (83), we can show that $\hat{B}^{(k-1)}_{L}$ and $\hat{B}^{(k-1)}_{R}$ satisfy Eq. (58). Furthermore, Eq. (59) follows from Eq. (76), which completes this proof. ∎